# Accelerating Distributed Optimization: A Primal-Dual Perspective on Local Steps

### v1

@@ Abstract
In distributed machine learning, efficient training across multiple agents with different data distributions poses significant challenges. Even with a centralized coordinator, current algorithms that achieve optimal communication complexity typically require either large minibatches or compromise on gradient complexity. In this work, we tackle both centralized and decentralized settings across strongly convex, convex, and nonconvex objectives. We first demonstrate that a basic primal-dual method, (Accelerated) Gradient Ascent Multiple Stochastic Gradient Descent (GA-MSGD), applied to the Lagrangian of distributed optimization inherently incorporates local updates, because the inner loops of running Stochastic Gradient Descent on the primal variable require no inter-agent communication. Notably, for strongly convex objectives, we show (Accelerated) GA-MSGD achieves linear convergence in communication rounds despite the Lagrangian being only linear in the dual variables. This is due to a unique structural property where the dual variable is confined to the span of the coupling matrix, rendering the dual problem strongly concave. When integrated with the Catalyst framework, our approach achieves nearly optimal communication complexity across various settings without the need for minibatches. Moreover, in stochastic decentralized problems, it attains communication complexities comparable to those in deterministic settings, improving over existing algorithms.
@@ 1 Introduction
The trend toward larger machine learning models and training with more extensive datasets has introduced new challenges in distributed training. The transmission of gradient or model information incurs substantial communication overhead not only due to bandwidth constraints but also because of latency and the need for privacy-preserving protocols [15]. Additionally, when each client possesses a substantial pool of local data, gradient estimates are usually noisy. These factors challenge traditional algorithms that perform communication after each local update with exact gradients [31, 45, 32, 59, 48]. A popular method known as FedAvg [28], or local SGD, entails performing multiple iterations of Stochastic Gradient Descent (SGD) locally before a single round of communication. However, this algorithm suffers from a suboptimal convergence rate and is impacted by the level of heterogeneity, when the local functions or data can vary significantly between clients [63, 11].
@@
In the centralized setup, Accelerated SGD with minibatches—whose size scales inversely with the target accuracy—can achieve optimal communication and sample complexities in both strongly convex and general convex settings [51]. However, employing such large minibatches is typically impractical in real-world applications. Although many algorithms that replace minibatches with local updates have been introduced [33, 38, 11], they struggle to achieve these optimal complexities. For example, the best-known communication complexity for local methods in convex settings remains at $\mathcal{O}(\epsilon^{-1})$ [11], compared to the optimal $\mathcal{O}(\epsilon^{-\frac{1}{2}})$ [50]. In strongly convex settings, with deterministic gradients, while local algorithms may reach the optimal communication complexity of $\widetilde{\mathcal{O}}(\sqrt{\kappa})$–where $\kappa$ denotes the condition number–they do so at the cost of suboptimal gradient complexity [29]; with noisy gradients, the optimal communication complexity is achieved only asymptotically [10]. Thus, designing local methods that can perform comparably to minibatch Accelerated SGD remains a challenging problem.
@@
In the decentralized setup, local Decentralized SGD [21, 12], a counterpart to FedAvg, still suffers from performance degradation due to data heterogeneity. Recently, several local methods have been introduced to improve this. For stochastic strongly convex problems, the best-known communication complexity is $\widetilde{\mathcal{O}}(\kappa^{2}(1-\sigma_{2})^{-1})$ with sufficient local steps [1], where $\sigma_{2}$ is the second largest singular value of the weight matrix. Yet, this complexity is still far from the optimal $\widetilde{\mathcal{O}}(\sqrt{\kappa/(1-\sigma_{2})})$ complexity observed in the deterministic setting [46, 42]. For stochastic convex problems, the best-known communication complexity is $\mathcal{O}((1-\sigma_{2})^{-1}\epsilon^{-1})$ [1], which starkly contrasts with the optimal $\mathcal{O}((1-\sigma_{2})^{-\frac{1}{2}}\epsilon^{-\frac{1}{2}})$ complexity seen in the deterministic setting [43, 20]. Consequently, it remains unclear whether communication complexities can be further improved in stochastic settings.
@@
(A) Centralized: When the objective is strongly convex, we show that GA-MSGD achieves linear convergence in the number of outer loops by establishing that the dual function of the Lagrangian is strongly concave. This is attributed to the coupling matrix between the primal and dual variables being full-rank. When integrated it with the Catalyst [23], we introduce a range of new results, including: (a) the first $\widetilde{\mathcal{O}}(\epsilon^{-\frac{1}{2}})$ communication complexity for convex settings without using minibatches; (b) the first local method that attains $\widetilde{\mathcal{O}}(\sqrt{\kappa})$ communication and gradient complexities simultaneously for deterministic strongly convex problems; (c) the first non-asymptotic $\widetilde{\mathcal{O}}(\sqrt{\kappa})$ communication complexity without minibatches for stochastic strongly convex problems. As a side result, we combine Catalyst with SCAFFOLD [11] to further attain linear speed-up [^1] In this work, we define an algorithm as having linear speedup if its sample complexity includes all terms associated with stochastic gradient variance, denoted as $\sigma^{2}$, divided by the total number of clients.. Results for strongly convex and convex settings are summarized in Table 1 and 2, respectively.
@@
(B) Decentralized: When the objective is strongly convex, we show that Acc-GA-MSGD converges linearly by proving that the dual function of the Lagrangian is strongly concave within the span of the coupling matrix between the primal and dual variables. When integrated with the Catalyst framework, we significantly improve communication efficiencies across all stochastic settings, provided that sufficient local steps are taken. Specifically, in both strongly convex and convex stochastic settings, our approach achieves communication complexities that match with the lower bounds in deterministic settings, up to logarithmic terms [42, 43]. As a side result, we provide a better communication complexity for LED [1] with our primal-dual perspective. We summarize the results in Table 3.
@@
In summary, our results demonstrate that with appropriate local methods, we can achieve communication complexities in stochastic settings that nearly match those in deterministic settings, provided that sufficient local steps are taken. Furthermore, a straightforward minimax optimization algorithm applied to the Lagrangian can serve as an effective local method for achieving such outcomes.
@@
Table 3: Communication Complexities for Decentralized Stochastic Problems. This table compares communication complexity when algorithms can use sufficient local samples for minibatches or local steps. Local-DSGD [12] needs additional bounded heterogeneity assumption. $\kappa$ represents the condition number, and $\sigma_{2}$ is the second largest singular value of the network weight matrix $W$. Stochastic GT [13] has another network-related constant $c$ with $c\geq 1-\sigma_{2}$.
@@
[b] Algorithms Strongly Convex Convex Nonconvex Local-DSGD [12] $\widetilde{\mathcal{O}}(\kappa(1-\sigma_{2})^{-1})$ $\mathcal{O}((1-\sigma_{2})^{-1}\epsilon^{-1}))$ $\mathcal{O}((1-\sigma_{2})^{-1}\epsilon^{-2})$ Stochastic GT [13] $\widetilde{\mathcal{O}}(\kappa(1-\sigma_{2})^{-1}c^{-1})$ $\mathcal{O}((1-\sigma_{2})^{-1}c^{-1}\epsilon^{-1})$ $\mathcal{O}((1-\sigma_{2})^{-1}c^{-1}\epsilon^{-2})$ K-GT [27] N/A N/A $\mathcal{O}((1-\sigma_{2})^{-2}\epsilon^{-2})$ LED [1] $\widetilde{\mathcal{O}}(\kappa^{2}(1-\sigma_{2})^{-1})$ $\mathcal{O}((1-\sigma_{2})^{-1}\epsilon^{-1}))$ $\mathcal{O}((1-\sigma_{2})^{-1}\epsilon^{-2}))$ LED $\widetilde{\mathcal{O}}(\kappa(1-\sigma_{2})^{-1})$ N/A N/A Acc-GA-MSGD $\widetilde{\mathcal{O}}(\kappa^{\frac{1}{2}}(1-\sigma_{2})^{-\frac{1}{2}})$ N/A N/A Catalyst+Acc-GA-MSGD $\widetilde{\mathcal{O}}(\kappa^{\frac{1}{2}}(1-\sigma_{2})^{-\frac{1}{2}})$ $\widetilde{\mathcal{O}}((1-\sigma_{2})^{-\frac{1}{2}}\epsilon^{-\frac{1}{2}}))$ $\widetilde{\mathcal{O}}((1-\sigma_{2})^{-\frac{1}{2}}\epsilon^{-2}))$ Lower bound [42, 43] $\widetilde{\Omega}(\kappa^{\frac{1}{2}}(1-\sigma_{2})^{-\frac{1}{2}})$ $\Omega((1-\sigma_{2})^{-\frac{1}{2}}\epsilon^{-\frac{1}{2}}))$ N/A
@@ 1.1 Related Work
In the decentralized setup, local methods, such as ProxSkip (also known as Scaffnew) [29] and DualFL [38], attain optimal communication complexity $\widetilde{\mathcal{O}}(\sqrt{\kappa})$ but suboptimal gradient complexity $\widetilde{\mathcal{O}}(\kappa)$ in the deterministic setting [29]. In the stochastic setting, the analysis of ProxSkip was recently tightened, resulting in a near-optimal communication complexity $\widetilde{\mathcal{O}}(\sqrt{\kappa})$ [10], though only asymptotically. Additionally, it still lacks linear speedup, and tuning the stepsizes requires solving an additional cubic equation. In the decentralized setup, the stochastic variant of Gradient Tracking with sufficiently large minibatches achieves the communication complexity of $\widetilde{\mathcal{O}}(\frac{\kappa}{(1-\sigma_{2})c})$ [13, 39], where $c\geq\Theta(1-\sigma_{2})$ is another network-related constant. Recently, Exact Diffusion with local steps [1] achieves a complexity of $\widetilde{\mathcal{O}}(\frac{\kappa^{2}}{1-\sigma_{2}})$.
@@
While minibatch Accelerated SGD achieves the optimal communication complexity of $\mathcal{O}(\epsilon^{-1/2})$ [51], the best performance for algorithms without minibatches remains at $\mathcal{O}(\epsilon^{-1})$, including SCAFFOLD [11] and RandProx [5]. In addition, decentralized algorithms can attain nearly optimal communication complexity of $\mathcal{O}((1-\sigma_{2})^{-\frac{1}{2}}\epsilon^{-\frac{1}{2}})$ in deterministic settings [20, 49]. However, the best-known stochastic decentralized algorithm only achieves a communication complexity of $\mathcal{O}((1-\sigma_{2})^{-1}\epsilon^{-1})$ [1].
@@
In the centralized setting, under bounded heterogeneity assumptions, local SGD attains a gradient complexity of $\mathcal{O}(\epsilon^{-4})$ and a communication complexity of $\mathcal{O}(\epsilon^{-3})$ [58]. Several algorithms, including [11, 4, 62, 1, 22], achieve near-optimal communication and gradient complexities without bounded heterogeneity. Some of them additionally achieve linear speedup [4]. In the decentralized setting, Gradient Tracking and Exact Diffusion, when combined with local steps, achieve communication complexities of $\mathcal{O}((1-\sigma_{2})^{-2}\epsilon^{-2})$ [27] and $\mathcal{O}((1-\sigma_{2})^{-1}\epsilon^{-2})$ [1], respectively.

### v2

@@ Abstract
In distributed machine learning, efficient training across multiple agents with different data distributions poses significant challenges. Even with a centralized coordinator, current algorithms that achieve optimal communication complexity typically require either large minibatches or compromise on gradient complexity. In this work, we tackle both centralized and decentralized settings across strongly convex, convex, and nonconvex objectives. We first demonstrate that a basic primal-dual method, (Accelerated) Gradient Ascent Multiple Stochastic Gradient Descent (GA-MSGD), applied to the Lagrangian of distributed optimization inherently incorporates local updates, because the inner loops of running Stochastic Gradient Descent on the primal variable require no inter-agent communication. Notably, for strongly convex objectives, (Accelerated) GA-MSGD achieves linear convergence in communication rounds despite the Lagrangian being only linear in the dual variables. This is due to a structural property where the dual variable is confined to the span of the coupling matrix, rendering the dual problem strongly concave. When integrated with the Catalyst framework, our approach achieves nearly optimal communication complexity across various settings without the need for minibatches.
@@ 1 Introduction
The trend toward larger machine learning models and training with more extensive datasets has introduced new challenges in distributed training. The transmission of gradient or model information incurs substantial communication overhead not only due to bandwidth constraints but also because of latency and the need for privacy-preserving protocols [24]. Additionally, when each client possesses a substantial pool of local data, gradient estimates are usually noisy. These factors challenge traditional algorithms that perform communication after each local update with exact gradients [43, 57, 45, 74, 60]. A popular method known as FedAvg [39], or local SGD, entails performing multiple iterations of Stochastic Gradient Descent (SGD) on local objectives between communications. However, this algorithm suffers from a suboptimal convergence rate and is impacted by the level of heterogeneity, when the local functions or data can vary significantly between clients [78, 19].
@@
In the centralized setup, Accelerated SGD with minibatches, whose size scales inversely with the target accuracy, achieve optimal communication and sample complexities in both strongly convex and general convex settings [64]. However, employing such large minibatches is typically impractical in real-world applications. Although many algorithms replace minibatches with local (stochastic) gradient steps—termed here as local methods—they typically do not achieve these optimal complexities [46, 51, 19]. In strongly convex settings, with deterministic gradients, while local algorithms reach the optimal communication complexity of $\widetilde{\mathcal{O}}(\sqrt{\kappa})$, where $\kappa$ denotes the condition number, they do so at the cost of suboptimal gradient complexity [40]; with noisy gradients, the optimal communication complexity is achieved only asymptotically [18]. Thus, designing local methods that can perform comparably to minibatch Accelerated SGD remains a challenging problem.
@@
In the decentralized setup, decentralized parallel SGD (D-PSGD) with minibatch [30], a stochastic variant of the subgradient method in [44], and decentralized local SGD [21], a decentralized counterpart to FedAvg, still suffer from performance degradation due to data heterogeneity and suboptimal convergence rates. Communication complexity lower bounds have been established at $\Omega(\kappa^{\frac{1}{2}}(1-\sigma_{2})^{-\frac{1}{2}})$ for strongly convex objectives and $\Omega((1-\sigma_{2})^{-\frac{1}{2}}\epsilon^{-2})$ for convex objectives [55, 56], where $\sigma_{2}$ is the second largest singular value of the weight matrix. Recently, stochastic algorithms with local steps achieved complexities of $\widetilde{\mathcal{O}}(\kappa^{2}(1-\sigma_{2})^{-1})$ and $\mathcal{O}((1-\sigma_{2})^{-1}\epsilon^{-1})$ for strongly convex and convex objectives, respectively [1], indicating a gap with the lower bounds. In the nonconvex setting, the optimal communication complexity is $\Omega((1-\sigma_{2})^{-\frac{1}{2}}\epsilon^{-2})$ and is achieved by a stochastic algorithm with minibatches [37]. However, a local method attains a suboptimal complexity of $\mathcal{O}((1-\sigma_{2})^{-1}\epsilon^{-2})$ [1]. This raises questions about whether local methods can reach optimal communication complexities in stochastic decentralized settings.
@@
(A) Centralized: When the objective is strongly convex, we show that GA-MSGD achieves linear convergence in the number of outer loops by establishing that the dual function of the Lagrangian is strongly concave. This is attributed to the coupling matrix between the primal and dual variables being full-rank. When integrated it with the Catalyst [32], notable results include: (a) the first local method that attains $\widetilde{\mathcal{O}}(\sqrt{\kappa})$ communication and gradient complexities simultaneously for deterministic strongly convex problems; (b) the first non-asymptotic $\widetilde{\mathcal{O}}(\sqrt{\kappa})$ communication complexity without minibatches for stochastic strongly convex problems. As a side result, we combine Catalyst with SCAFFOLD [19] to further attain linear speed-up [^1] In this work, we define an algorithm as having linear speedup if its sample complexity includes all terms associated with stochastic gradient variance, denoted as $\sigma^{2}$, divided by the total number of clients.. Results for strongly convex and convex settings are summarized in Table 1 and 2, respectively.
@@
(B) Decentralized: When the objective is strongly convex, we demonstrate that Acc-GA-MSGD achieves linear convergence by observing that the dual function of the Lagrangian is strongly concave within the span of the coupling matrix between the primal and dual variables. By integrating Acc-GA-MSGD with the Catalyst framework, we achieve near-optimal communication efficiencies across strongly convex, convex, and nonconvex settings, provided that sufficient local steps are taken. This matches the lower bounds established in [55, 56, 37]. The results are summarized in Table 3.
@@
In summary, our results reveal that by employing suitable local methods, we can achieve the near-optimal communication complexities in stochastic settings, without the need for large minibatches. Furthermore, a straightforward minimax optimization algorithm applied to the Lagrangian can serve as an effective local method for achieving such outcomes. While the algorithms resemble the (accelerated) dual ascent methods in [55, 61], direct access to the gradient of the conjugate function is not available in our work.
@@
Table 3: Communication Complexities for Decentralized Stochastic Problems. This table compares communication complexity when algorithms can use sufficient local samples for minibatches or local steps. assumes a heterogeneity measure is bounded by $\varsigma$. Δminibatch is needed for the best communication complexity. $\kappa$ represents the condition number, and $p\triangleq 1-\sigma_{2}$, where $\sigma_{2}$ is the second largest singular value of the network weight matrix $W$. $M$ is the number of clients. $C_{0}$ in D-MSAG is specified in [12]. Stochastic GT [22] has another network-related constant $c$ with $c\geq 1-\sigma_{2}$.
@@
[b] Algorithms Strongly Convex Convex Nonconvex Local-DSGD [21] $\widetilde{\mathcal{O}}(\kappa p^{-1}+\sqrt{\kappa}\varsigma p^{-1}\epsilon^{-\frac{1}{2}})$ $\mathcal{O}(p^{-1}\epsilon^{-1}+\varsigma p^{-1}\epsilon^{-\frac{3}{2}})$ $\mathcal{O}(p^{-1}\epsilon^{-2}+\varsigma p^{-1}\epsilon^{-3})$ D-PSGD [30]Δ N/A N/A $\mathcal{O}(\varsigma Mp^{-2}c^{-1}\epsilon^{-2})$ DeTAG [37]Δ N/A N/A $\widetilde{\mathcal{O}}(p^{-\frac{1}{2}}\epsilon^{-2})$ D-MSAG [12]Δ $\widetilde{\mathcal{O}}(\kappa^{\frac{1}{2}}p^{-\frac{1}{2}}+C_{0}\sqrt{\kappa}p^{-\frac{1}{2}}(M\epsilon)^{-\frac{1}{4}})$ N/A N/A Stochastic GT [22]Δ $\widetilde{\mathcal{O}}(\kappa p^{-1}c^{-1})$ $\mathcal{O}(p^{-1}c^{-1}\epsilon^{-1})$ $\mathcal{O}(p^{-1}c^{-1}\epsilon^{-2})$ K-GT [36] N/A N/A $\mathcal{O}(p^{-2}\epsilon^{-2})$ LED [1] $\widetilde{\mathcal{O}}(\kappa^{2}p^{-1})$ $\mathcal{O}(p^{-1}\epsilon^{-1}))$ $\mathcal{O}(p^{-1}\epsilon^{-2})$ Acc-GA-MSGD $\widetilde{\mathcal{O}}(\kappa^{\frac{1}{2}}p^{-\frac{1}{2}})$ N/A N/A Catalyst+Acc-GA-MSGD $\widetilde{\mathcal{O}}(\kappa^{\frac{1}{2}}p^{-\frac{1}{2}})$ $\widetilde{\mathcal{O}}(p^{-\frac{1}{2}}\epsilon^{-\frac{1}{2}})$ $\widetilde{\mathcal{O}}(p^{-\frac{1}{2}}\epsilon^{-2})$ Lower bound [55, 56, 37] $\widetilde{\Omega}(\kappa^{\frac{1}{2}}p^{-\frac{1}{2}})$ $\Omega(p^{-\frac{1}{2}}\epsilon^{-\frac{1}{2}})$ $\Omega(p^{-\frac{1}{2}}\epsilon^{-2})$
@@ 1.1 Related Work
In the centralized setup, local methods such as ProxSkip (also known as Scaffnew) [40] and DualFL [51] attain optimal communication complexity $\widetilde{\mathcal{O}}(\sqrt{\kappa})$ but suboptimal gradient complexity $\widetilde{\mathcal{O}}(\kappa)$ in the deterministic setting [40]. In the stochastic setting, the communication complexity of ProxSkip was recently tightened to the near-optimal $\widetilde{\mathcal{O}}(\sqrt{\kappa})$, though only asymptotically [18]. In the decentralized setup, deterministic algorithms have achieved near-optimal communication complexity $\widetilde{\Omega}(\kappa^{\frac{1}{2}}(1-\sigma_{2})^{-\frac{1}{2}})$ [55, 62, 29, 28, 58]. Notably, Uribe et al. [62] incorporated local steps in an algorithm similar to Scaffnew. Fallah et al. [12] considered decentralized multi-stage Accelerated Stochastic Gradient (D-MASG) that can converge to the exact solution in the stochastic setting. The stochastic variant of Gradient Tracking [52] with sufficiently large minibatches achieves a communication complexity of $\widetilde{\mathcal{O}}(\frac{\kappa}{(1-\sigma_{2})c})$, where $c\geq\Theta(1-\sigma_{2})$ is a network-related constant [22]. Recently, Exact Diffusion with local steps achieved a complexity of $\widetilde{\mathcal{O}}(\frac{\kappa^{2}}{1-\sigma_{2}})$ [1].
@@
In the centralized setup, while minibatch Accelerated SGD achieves the optimal communication complexity of $\mathcal{O}(\epsilon^{-1/2})$ [64], many federated learning algorithms that do not require minibatches, including SCAFFOLD [19] and RandProx [9], remain at $\mathcal{O}(\epsilon^{-1})$. Additionally, decentralized algorithms can attain the optimal communication complexity of $\mathcal{O}((1-\sigma_{2})^{-\frac{1}{2}}\epsilon^{-\frac{1}{2}})$ in deterministic settings [29, 62, 56, 28], with some exploring algorithms with local steps [62]. A stochastic algorithm with local steps achieves a communication complexity of $\mathcal{O}((1-\sigma_{2})^{-1}\epsilon^{-1})$ [1].
@@
In the centralized setting, under bounded heterogeneity assumptions, local SGD attains a gradient complexity of $\mathcal{O}(\epsilon^{-4})$ and a suboptimal communication complexity of $\mathcal{O}(\epsilon^{-3})$ [73]. Several algorithms, including [19, 8, 77, 1, 31], achieve near-optimal communication and gradient complexities without bounded heterogeneity. Some of them additionally achieve linear speedup [8]. In the decentralized setting, local SGD suffers from suboptimal communication complexity in $\epsilon$ [21], while decentralized minibatch SGD achieves $\mathcal{O}((1-\sigma_{2})^{-2}\epsilon^{-2})$ with bounded heterogeneity [30]. It was further improved with techniques such as PushSum [2] and gradient tracking [67]. The lower bound of $\Omega((1-\sigma_{2})^{-\frac{1}{2}}\epsilon^{-2}))$ communication complexity was established in [37], and a nearly matched algorithm was also provided, but requires a large batch size. Gradient Tracking and Exact Diffusion, when combined with local steps, achieve communication complexities of $\mathcal{O}((1-\sigma_{2})^{-2}\epsilon^{-2})$ [36] and $\mathcal{O}((1-\sigma_{2})^{-1}\epsilon^{-2})$ [1], respectively.
(Primal)-Dual Algorithms:
There is a long history of using primal-dual algorithms to solve the (augmented) Lagrangian. The Alternating Direction Method of Multipliers (ADMM) is a popular technique for solving the augmented Lagrangian [6, 63, 38]. Several works have suggested using linearized primal subproblems [7, 16, 34, 3], with a closed form for the primal update that involves a one-step (stochastic) gradient descent in the primal. Many algorithms, including EXTRA [57, 41, 28] and DIGing [45], are interpreted or motivated by the Gradient Descent Ascent algorithm on the augmented Lagrangian. The quadratic penalty term related to the gossip matrix in the augmented Lagrangian prevents a straightforward extension of these methods to include multiple gradient descent steps in the primal variable without communication. These methods also closely resemble linearized ADMM [16]. Mokhtari and Ribeiro [41] showed linear convergence for stochastic strongly convex problems when each agent has finite-sum objectives by incorporating incremental averaging gradients used in variance-reduction literature. Lan et al. [27] considers a primal-dual method on the non-augmented Lagrangian for non-smooth convex problems. The most closely related works to ours are [61, 13, 55, 62], which executed (accelerated) dual ascent in the dual problem. There are two main differences between our work with the first three works: they assume the exact dual gradients are available, whereas we use stochastic gradient descent in the primal function to find inexact dual gradients; and their convergence is explicitly provided only for the dual variable, rather than the primal variable. Specifically, linear convergence is provided in [55], but the proof uses the Hessian of the conjugate function, whose existence requires additional assumptions, such as twice differentiability of the original objective. Uribe et al. [62] used Accelerated Gradient Descent to approximate the dual gradient but did not address stochastic or non-convex settings.
@@ 2 Preliminary
Li and Lin [28] extended the Catalyst framework to decentralized optimization by demonstrating that applying Catalyst individually to each client essentially equates to applying the original Catalyst to the average of the clients’ variables (see Section C.3 for more details). This adaptation allows us to apply Theorem 1 in decentralized settings as well.